what is the equation of the line that has a slope of 6/7 and a point of (2,-8)?
Answer by FrankM(1040)

Point slope form
(y-y1)=m(x-x1)
y-(-8)=6/7(x-2)
y+8=6/7(x-2)
Let's multiply both sides by 7
7y+56=6(x-2)
7y+56=6x-12
-6x+7y=-68 or
6x-7y=68 general form